# captionUnknown document class

### v2

@@ Abstract
This paper extends the fall prediction algorithm from [1] to a real-time/online setting, implemented in both hardware and simulation. This yields results comparable to the offline version, maintaining a zero false positive rate, sufficient lead time, and accurate lead time prediction. Additionally, it achieves a high recovery rate. The paper also evaluates the fall prediction algorithm against omnidirectional faults and introduces an improved algorithm capable of reliably predicting falls and lead times across a wider range of faults in full-sized robots. Compared to [1], the proposed algorithm performs significantly better across all metrics, such as false positive rate, lead time, accuracy, and response time, demonstrating the algorithm’s efficacy for real-time fall prediction in bipedal robots.
@@ I Introduction
Humanoid robots have attracted significant attention over the past year, with a surge in companies developing general-purpose humanoids for both home and factory applications. Although some humanoids have already been deployed in home and factory environments, their widespread adaptation, especially in unstructured or unknown environments, remains limited [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. A key challenge in broadly deploying humanoid robots in these environments is their high-dimensional, hybrid nature, which, paired with their smaller support polygons, makes it particularly difficult to maintain stable motion, especially in unstructured terrains in the presence of disturbances and uncertainties. While a well-designed controller can mitigate the effects of disturbances and uncertainties, the unpredictable nature of real-world operations makes falls inevitable. As a result, it is imperative to implement fall prediction algorithms alongside robust controllers, especially in real-time operations. Therefore, this paper extends the previously proposed fall prediction framework [1] to real-time operations, analyzes the algorithm’s robustness to omnidirectional disturbances during the task of standing, and offers pathways to expand its capabilities. Note that the task of standing was chosen in [1] as a precursor to scaling up to more dynamic motions.
@@ I-A Background
The overarching objective of a fall prediction algorithm is to accurately predict falls with sufficient time to implement a recovery strategy. As discussed in [1] and [12], faults can be a precursor to a bipedal robot falling. Faults are defined as unforeseen deviations in one or more operational variables. Based on their temporal behavior, they can be categorized as either abrupt, incipient, or intermittent. Abrupt faults are sudden, incipient faults develop gradually, and intermittent faults occur sporadically. Intermittent faults are comprised of abrupt and incipient faults [13]. In this paper, we refer to faults that lead to a fall as critical faults. The term lead time is used to describe the difference in time between the prediction of a fall and its actual occurrence, while response time is the difference in time between when a critical fault is introduced and when a fall is predicted. A trajectory or state is considered unsafe if it leads to a fall, and safe if it does not.
@@ I-B Literature Review
While there are numerous fall prediction algorithms for bipedal robots, ranging from static threshold-based methods to machine learning approaches [14, 15], the implementation of algorithms capable of real-time/online prediction of falls caused by abrupt, incipient and intermittent faults in full-sized humanoids, in hardware, remains limited.
For clarity, in addition to the detection of the three types of faults and the size of the robot, we also distinguish between four experimental settings: offline with simulation data, offline with hardware data, online in simulation, and online in hardware.
@@
The fall prediction algorithm proposed in [16] uses sensor fusion to predict falls in real-time, across various terrains and during dynamic motions such as squatting, in simulation and hardware for the 1.92m WALK-MAN humanoid [17]. However, the analysis is limited to falls caused by abrupt faults. Similarly, the SVM-based algorithm from [18] is tested online in hardware and simulation for the 1.6m BHR-6 robot during walking and standing tasks, and only considers abrupt faults.
@@
In contrast, while [19], [20], [21], and [22] do not explicitly address all three faults, their analyses are based on scenarios that may encompass all temporal fault categories—for instance, varying the robot’s initial state [19], robot-to-robot interactions [20], walking on uneven terrain [21], and motor failures [22]. Despite the potential inclusion of all three faults, the analyses in [19] and [21] are limited to offline simulations of the 1.43m planar bipedal robot Rabbit [23] and the 90cm NUgus humanoid robot, respectively. While [22] uses both hardware and simulation data to evaluate their algorithm on the 1.82m Q1 humanoid robot and a Cassie bipedal robot—originally 1m tall[24]—fitted with a torso, their analysis does not include online hardware experiments. On the other hand, [20] conducts their analysis online in both hardware and simulation, though it is limited to the 0.57m tall Nao H25 robot [25, 26].
In comparison to the other approaches, building on previous work in [12], [1] develops a 1D CNN model capable of predicting falls from all three faults and estimating lead times for the 1.6m tall bipedal robot Digit [27], both in simulation and hardware. The algorithm was validated on a new open-source dataset for the Digit platform [28]. However, the study is limited to offline prediction in a planar standing task, where faults are simulated by applying forces to the robot’s back center. Despite this limitation, the model demonstrates high prediction accuracy in offline settings.
@@ I-C Objective
Therefore, the objective of this paper is to extend [1]’s algorithm, which we term the nominal algorithm, to an online setting, both in hardware and simulation, and to improve the algorithm’s performance to omnidirectional faults during the task of standing, while adhering to the requirements outlined in [1]. Specifically, this paper aims to:
  • Maintain the nominal algorithm’s goal of reliably detecting imminent falls caused by abrupt, incipient, and intermittent faults, with sufficient lead time for a full-sized bipedal robot, while also accurately predicting the lead time.
  • Implement the nominal fall prediction algorithm online, both in hardware and in simulation.
  • Evaluate the performance of the nominal algorithm on critical faults applied in the sagittal and frontal plane at various angles and locations on the robot, using the same metrics as [1]: false positive rates, lead time, and response time.
  • Propose a method for robustifying the nominal algorithm against a broader set of faults.
As in [1], the robot is considered to have fallen if its center of mass is less than 10 percent of its initial height, and the minimum desired lead time is set to 0.2 seconds.
Achieving the objective outlined above is challenging due to several factors: the masking effects of the controller, the sporadic nature of intermittent faults, the trade-off between lead time and false positive rates, and the crowding effect of incipient faults [29]. Furthermore, the decreasing number of available data points as the lead time increases [1], along with the diverse range of faults, complicates the prediction of falls. Implementing the algorithm online introduces additional challenges, including the need for fast inference times and the management of communication delays between different components of the fall prediction algorithm and other processes, such as the controller.
@@ I-D Contributions
  • Extension of the open-source dataset [28] to include critical and non-critical faults, introduced by applying angled and non-angled forces at various locations in the sagittal and frontal planes of the robot
  • Analysis of the nominal fall prediction algorithm against a broader range of faults.
  • Introduction of a fall prediction algorithm capable of predicting falls caused by a wider range of critical faults while also predicting lead time.
II System Overview: Nominal Fall Prediction Algorithm, Hardware Description, and Experimental Setup
This section provides a brief overview of the fall prediction algorithm introduced in [1], along with a concise summary of the hardware and experimental setup used in this paper, which are similar to [1].
II-A Fall Prediction Algorithm Framework
The 1D CNN-based fall prediction algorithm introduced in [1] consists of three parts: a critical fault classifier, a lead time classifier, and a lead time regressor. The critical fault classifier identifies critical faults, while the lead time classifier and the regressor work together to predict the lead time. The lead time classifier divides the lead time into three intervals: $[0,1]$, $(1,2]$, and $(2,H]$. If the lead time is classified into the $[0,1]$ interval, the lead time regressor calculates the lead time. Otherwise, the lead time is considered as the infimum of the interval. To prioritize the most recent data points, sliding windows with 10 data points are used. The algorithm is trained with trajectories sampled at approximately 40 Hz. Figure 1 depicts the algorithm’s framework.
Figure 1: The fall prediction algorithm proposed in [1] is comprised of three components: a critical fault classifier, a lead time classifier, and a lead time regressor. The green boxes depict the algorithm’s predicted lead time.
@@ II-B Hardware Description
We utilize the bipedal robot Digit [27] developed by Agility Robotics. Digit is approximately 1.6m, weighs 48 kg and has 30 degrees of freedom, 20 actuated joints, and an integrated perception system. Figure 2 illustrates Digit’s kinematic architecture.
@@
Figure 2: Kinematics architecture of the Digit robot by Agility Robotics [27, 1].
II-C Experimental Setup
II-C1 Simulation
We employ the use of Agility’s MuJoCo-based simulator with a standing controller designed to maintain the center of mass and zero moment point inside the support polygon [30]. This controller will be referred to as the nominal controller. The nominal controller can stabilize incipient faults up to a force of 28.8N and abrupt faults up to 207.4N [1]. Abrupt and incipient faults are emulated in simulation by applying impulsive and trapezoidal pushing forces, respectively, to the back center of the robot, denoted as region F in Figure 5. To introduce slight oscillations to the robot’s standing posture, random abrupt forces within the range of 0 - 202.4N are introduced, but only the windows during and after the fault introduction are passed to the fall prediction algorithm.
II-C2 Hardware
In contrast to the simulation setup, where faults are applied to Digit’s back center, this location was avoided during hardware experiments as it houses the onboard computer and is vulnerable to damage. Instead, faults are introduced in region C on the back of the torso, as shown in Figure 5, by applying pushing forces with a pole. To capture the gradual nature of incipient faults, the pole is first positioned on the robot before applying force. To prevent potential damage during experiments, the robot is secured to a gantry using a slack cable, and motor power is disabled upon detecting a fall.
III Online Fall Prediction
@@
This section assesses the performance of the nominal fall prediction algorithm when executed online in both simulation and hardware.
III-A Experimental Setup
In the hardware setup, we do not have direct access to the main computer, so we interact with the robot by sending commands and receiving state information over UDP at approximately 1 kHz. This information is processed by the nominal controller, which publishes topics via ROS also at roughly 1 kHz. The robot is connected to an Alienware m15 computer with an Intel i9-12900H processor via Ethernet. This computer runs both the fall prediction algorithm and the nominal controller. The simulation emulates this asynchronous setup. However, it is run on a machine with an Intel i9-7900x processor.
@@ III-A Experimental Setup
To run the fall prediction algorithm online, the robot is first initialized in a standing position with the nominal controller. The robot’s state and the torques commanded by the nominal controller are then queried via ROS [31] at a rate of 400Hz. The fall prediction algorithm, however, runs at 40 Hz. Therefore, features for the algorithm are computed using newly queried data every 40 Hz. Once enough features are available to form a window, model inference is performed continuously until a fall is predicted. Once a fall is predicted, the recovery controller, if one is available, is activated. Note 40 Hz is chosen to match the sampling frequency of the trajectories in [1].
@@
Building on the offline analysis presented in [1], the algorithm’s performance is evaluated using lead time, false positive rate, and response time. In addition to these metrics, we introduce a new metric to assess the algorithm’s effectiveness: the recovery rate. The recovery rate measures the number of unsafe trajectories that are stabilized by switching to a recovery controller when a critical fault is detected. We report the recovery rate only for simulation tests due to the difficulty of replicating the same fault multiple times in hardware. Since implementing a recovery controller falls outside the scope of this work, we utilize Agility’s base controller, which already incorporates a fall prevention strategy.
As in the offline tests, faults are emulated by applying horizontal forces of varying magnitudes to regions C and F in Figure 5 for hardware and simulation, respectively. The online evaluations are also limited to abrupt and incipient faults, as the algorithm’s effectiveness in detecting intermittent faults, as demonstrated in [1], depends on accurately identifying critical abrupt and incipient faults.
III-B Simulation Results
@@
To assess the fall prediction algorithm’s performance online in simulation and compare it to the offline results presented in [1], we replicate the 200 abrupt and incipient tests used in the offline evaluation and apply the algorithm to these tests as described in Section III-A. The average runtime of the fall prediction algorithm is approximately 0.0045s, with the total inference time of the 1D CNN models being 0.004s and the computation of the features taking 0.0005s.
@@
As shown in Table I, the true average lead time for the online execution is 0.01s higher than the offline execution. This minor discrepancy in fault detection is primarily due to communication delays, which cause the online algorithm to occasionally run slower, averaging 40 Hz but sometimes dipping to lower rates. As a result, windows with features are sampled at intervals slightly greater or less than 40 Hz, leading to differences in the timing of fault detection. These timing variations also contribute to small changes in the average response time and predicted lead time, which decrease by 0.02s and increase by 0.04s, respectively, compared to the offline results. The difference between the average true lead time and predicted lead time is 0.03s for the online execution and 0.02s for the offline execution. Both the online and offline evaluations report an average false positive rate of 0. Given the small differences between the online and offline results, we can conclude that the algorithm’s online performance is closely aligned with the offline results. Furthermore, the 0.97 recovery rate for the unsafe trajectories demonstrates the effectiveness of running the fall prediction algorithm online. Figure 3 showcases the results of switching vs not switching to a recovery controller when the nominal fall prediction algorithm detects a critical fault.
@@
TABLE I: Online vs Offline Average Simulation Results of the Fall Prediction Algorithm Evaluated on Abrupt and Incipient Faults
Operation
Predicted
Lead
Time
True
Lead
Time
False
Positive
Rate
Response
Time
Recovery
Rate
Offline 1.18 1.16 0.0 0.44 N/A
@@
III-C Hardware Results
Given that [1] demonstrates the effectiveness of the fall prediction algorithm in detecting critical faults offline in both hardware and simulation, and as shown in the previous section, the algorithm’s offline performance is indicative of its online performance in simulation, we now seek to evaluate the algorithm’s inference time in hardware. To accomplish this, we execute the algorithm online in hardware 13 times, emulating 8 abrupt faults and 6 incipient faults in a manner similar to the setup described in [1]. Due to the challenges of replicating the same fault twice in hardware and to prevent potential damage to the robot, all tests are conducted with the recovery controller in the loop.
Out of the 8 abrupt faults, 4 are detected as critical faults, and 3 of the 6 incipient faults are identified as critical faults. The average predicted lead time of the entire algorithm is 1.04s. This is the same as the offline hardware results reported in [1] and 0.1s lower than the results obtained in the online simulation. This deviation can be attributed to the smaller number of hardware online tests, which do not encompass as broad a range of critical faults. Note that we only report the average predicted lead time, as measuring the force application time and hence the response time online without force sensors proved challenging. Additionally, false-positive rates are unknown since the recovery controller is triggered whenever falls are detected.
The average runtime of the fall prediction algorithm for the hardware setup is 0.0061s, which is comparable to the algorithm’s runtime in simulation. The 0.0016s difference in runtime is negligible, given the algorithm’s sampling rate of 0.025s (40Hz). Furthermore, the nominal controller publishes at about 1kHz in both hardware and simulation. The negligible difference in runtime for the fall prediction algorithm between simulation and hardware, the small sim-to-real gap demonstrated offline in [1], and the similarity in performance when executed offline and online in simulation, as shown in Table I, suggest that the algorithm’s performance in simulation reliably translates to hardware, further validating its real-world applicability. Figure 4 demonstrates the fall prediction algorithm detecting a critical fault online in hardware.

### v3

# Standing Tall: Sim to Real Fall Classification and Lead Time Prediction for Bipedal Robots

@@ Abstract
This paper extends a previously proposed fall prediction algorithm to a real-time (online) setting, with implementations in both hardware and simulation. The system is validated on the full-sized bipedal robot Digit, where the real-time version achieves performance comparable to the offline implementation while maintaining a zero false positive rate, an average lead time (defined as the difference between the true and predicted fall time) of 1.1s (well above the required minimum of 0.2s), and a maximum lead time error of just 0.03s. It also achieves a high recovery rate of 0.97, demonstrating its effectiveness in real-world deployment. In addition to the real-time implementation, this work identifies key limitations of the original algorithm, particularly under omnidirectional faults, and introduces a fine-tuned strategy to improve robustness. The enhanced algorithm shows measurable improvements across all evaluated metrics, including a 0.05 reduction in average false positive rate and a 1.19s decrease in the maximum error of the average predicted lead time.
@@ I Introduction
Humanoid robots have attracted significant attention over the past year, with a surge in companies developing general-purpose humanoids for home and factory applications. Although some humanoids have already been deployed in home and factory environments [ref:morgan_stanley_report] and their performance in the recent World Humanoid Robot Games in Beijing showcased their increased ability to handle complex situations[ref:wall_street_journal], their widespread adaptation, especially in unstructured or unknown environments, remains limited. A key challenge in broadly deploying humanoid robots in these environments is their high-dimensional, hybrid nature, which, paired with their smaller support polygons, makes it particularly difficult to maintain stable motion, especially in unstructured terrains in the presence of disturbances and uncertainties. Although a well-designed controller can mitigate the effects of disturbances and uncertainties, the unpredictable nature of real-world operations makes falls inevitable. This is especially critical in applications involving human-robot interaction where safety is paramount. As a result, it is imperative to implement fall prediction algorithms alongside robust controllers. Therefore, this paper extends the previously proposed fall prediction framework [ref:mungai2024fall] to real-time operations, analyzes the algorithm’s robustness to omnidirectional disturbances during the task of standing, and offers pathways to expand its capabilities. Standing was chosen in [ref:mungai2024fall] as a precursor to more dynamic behaviors.
@@ I-A Background
The overarching objective of a fall prediction algorithm is to accurately predict falls with sufficient time to implement a recovery strategy. As discussed in [ref:mungai2024fall] and [ref:mungai2023optimizing], faults can be a precursor to a bipedal robot falling. Faults are defined as unforeseen deviations in one or more operational variables. Based on their temporal behavior, they can be categorized as either abrupt, incipient, or intermittent. Abrupt faults are sudden, incipient faults develop gradually, and intermittent faults occur sporadically. Intermittent faults are comprised of abrupt and incipient faults [ref:ISERMANN19827]. In this paper, we refer to faults that lead to a fall as critical faults. The term lead time is used to describe the difference in time between the prediction of a fall and its actual occurrence, while response time is the difference in time between when a critical fault is introduced and when a fall is predicted. A trajectory or state is considered unsafe if it leads to a fall, and safe if it does not.
Figure 1: Real-time detection of a critical fault introduced between the first and second panels.
@@ I-B Literature Review
While numerous fall prediction algorithms have been proposed for bipedal robots—ranging from static threshold-based methods to machine learning approaches [ref:subburaman2023survey, ref:new_zhang2024fall]—the implementation of algorithms capable of real-time prediction of falls caused by abrupt, incipient, or intermittent faults in full-sized humanoid robots remains limited. Moreover, the prediction of lead time is also limited. Given the trade-off between false positive rate and lead time, and the dependence of the minimum required lead time on both the recovery algorithm and hardware platform, lead time prediction can significantly inform the selection and timing of an appropriate recovery strategy.
Fall prediction algorithms in the literature can be categorized based on several factors: the type of faults they address (abrupt, incipient, or intermittent), the size of the robot used, the experimental setting (offline with simulation data, offline with hardware data, online in simulation, or online in hardware), and whether they estimate lead time. In this section, we briefly go over the fall prediction algorithms in the literature based on these categorizations.
@@
The fall prediction algorithm proposed in [ref:subburaman] uses sensor fusion to predict falls in real-time, across various terrains and during dynamic motions such as squatting, in simulation and hardware for the 1.92m WALK-MAN humanoid [ref:walkman]. However, the analysis is limited to falls caused by abrupt faults. Similarly, the SVM-based algorithm from [ref:wu2021falling] is tested online in hardware and simulation for the 1.6m BHR-6 robot during walking and standing tasks, and only considers abrupt faults.
@@
In contrast, while [ref:new_igual2024time], [ref:tran], [ref:khaled_2025], and [ref:ZHANG2025104995] do not explicitly address all three faults, their analyses are based on scenarios that may encompass all temporal fault categories—for instance, varying the robot’s initial state [ref:new_igual2024time], robot-to-robot interactions [ref:tran], walking on uneven terrain [ref:khaled_2025], and motor failures [ref:ZHANG2025104995]. Despite the potential inclusion of all three faults, the analyses in [ref:new_igual2024time] and [ref:khaled_2025] are limited to offline simulations of the 1.43m planar bipedal robot Rabbit [ref:chevallereau2003rabbit] and the 90cm NUgus humanoid robot, respectively. While [ref:tran] conduct their analysis online in both hardware and simulation, it is limited to the 0.57m tall Nao H25 robot [ref:nao, ref:naoh25]. On the other hand, [ref:ZHANG2025104995] uses both hardware and simulation data to evaluate their algorithm on the 1.82m Q1 humanoid robot and a Cassie bipedal robot—originally 1m tall[ref:gong2019feedback]—fitted with a torso, however, their analysis does not include online hardware experiments.
[ref:mungai2024fall] proposes a 1D CNN model capable of predicting falls caused by abrupt, incipient, and intermittent faults, both in simulation and hardware, using the 1.6m tall bipedal robot, Digit [ref:ar_digit]. While comparable to [ref:ZHANG2025104995], the algorithm in [ref:mungai2024fall] is, to the best of our knowledge, the only one that also estimates lead time on a full-sized humanoid. Additionally, the authors provide an open-source dataset for reproducibility [ref:digit_dataset]. However, the study is limited to offline prediction in a planar standing task, where faults are simulated by applying external forces to the robot’s back.
@@ I-C Objective
Given the high offline accuracy in critical fault detection and lead time prediction of the fall prediction algorithm proposed by [ref:mungai2024fall], along with the availability of its dataset, the objective of this paper is to extend the algorithm to a real-time (online) setting and validate its performance in both hardware and simulation on the full-sized bipedal robot Digit. Furthermore, as the algorithm is only trained and evaluated on faults introduced by horizontal forces applied to the center back of the robot’s torso, this work seeks to identify its limitations under omnidirectional faults during a standing task and propose a fine-tuned strategy to improve robustness with minimal additional data. Note that the task of standing is chosen as a precursor to more dynamic motions. The algorithm proposed in [ref:mungai2024fall] is referred to as the nominal algorithm.
Specifically, this paper:
  • Maintains the nominal algorithm’s goal of reliably detecting imminent falls caused by abrupt, incipient, and intermittent faults, with sufficient lead time for a full-sized bipedal robot, while also accurately predicting lead time.
  • Adopts the nominal fall prediction algorithm to an online setting, both in hardware and in simulation.
  • Evaluates the performance of the nominal algorithm on critical faults applied in the sagittal and frontal plane at various angles and locations on the robot, using the same metrics: false positive rates, lead time, and response time.
  • Proposes a strategy for robustifying the nominal algorithm against a broader set of faults.
As in [ref:mungai2024fall], the robot is considered to have fallen if its center of mass is less than 10 percent of its initial height, and the minimum desired lead time is set to 0.2s.
Achieving the objective outlined above is challenging due to several factors: the masking effects of the controller, the sporadic nature of intermittent faults, the trade-off between lead time and false positive rates, and the crowding effect of incipient faults [ref:safaeipour2021survey]. Furthermore, the decreasing number of available data points as the lead time increases [ref:mungai2024fall], along with the diverse range of faults, complicates the prediction of falls. Implementing the algorithm online introduces additional challenges, including the need for fast inference times and the management of communication delays between different components of the fall prediction algorithm and other processes, such as the controller.
@@ I-D Contributions
  • Extension of the open-source dataset [ref:digit_dataset] to include critical and non-critical faults, introduced by applying angled and non-angled forces at various locations in the sagittal and frontal planes of the robot
  • Analysis and robustification of the nominal fall prediction algorithm against a broader range of faults.
While our study focuses on the nominal fall prediction algorithm for the bipedal robot Digit during standing, the techniques used are applicable to other algorithms, platforms, and tasks.
II Method: Nominal Fall Prediction Algorithm, Hardware Description
This section provides a brief overview of the nominal fall prediction algorithm, along with a description of the robot hardware used.
II-A Overview of Nominal Fall Prediction Algorithm Framework
The 1D CNN-based nominal fall prediction algorithm consists of three parts: a critical fault classifier, a lead time classifier, and a lead time regressor. The critical fault classifier identifies critical faults, while the lead time classifier and the regressor work together to predict the lead time. The lead time classifier divides the lead time into three intervals: $[0,1]$, $(1,2]$, and $(2,H]$. If the lead time is classified into the $[0,1]$ interval, the lead time regressor calculates the lead time. Otherwise, the lead time is considered as the infimum of the interval. To prioritize the most recent data points, sliding windows with 10 data points are used. The algorithm is trained with trajectories sampled at approximately 40 Hz. Fig. 2 depicts the algorithm’s framework.
Figure 2: The nominal fall prediction algorithm is comprised of three components: a critical fault classifier, a lead time classifier, and a lead time regressor. The green boxes depict the algorithm’s predicted lead time.
@@ II-B Hardware Description
We utilize the bipedal robot Digit [ref:ar_digit] developed by Agility Robotics. Digit is approximately 1.6m, weighs 48 kg and has 30 degrees of freedom, 20 actuated joints, and an integrated perception system. Fig. 3 illustrates Digit’s kinematic architecture.
@@
Figure 3: Kinematics architecture of the Digit robot by Agility Robotics [ref:ar_digit, ref:mungai2024fall].
III Online Implementation and Evaluation of Nominal Fall Prediction Algorithm
@@
This section presents the implementation of the nominal fall prediction algorithm online in both simulation and hardware and assesses it’s performance.
III-A Online Implementation
@@
As in the offline implementation, we employ the use of Agility’s MuJoCo-based simulator with a standing controller designed to maintain the center of mass and zero moment point inside the support polygon [ref:grant]. This standing controller will be referred to as the nominal controller. To run the fall prediction algorithm online, the robot is first initialized in a standing position with the nominal controller. The robot’s state and the torques commanded by the nominal controller are then queried via ROS [ref:ros] at a rate of 400Hz. The nominal fall prediction algorithm, however, runs at 40 Hz. Therefore, features for the algorithm are computed using newly queried data every 40 Hz. Once enough features are available to form a window, model inference is performed continuously until a fall is predicted. Once a fall is predicted, the recovery controller, if one is available, is activated. The average runtime of the nominal fall prediction algorithm is approximately 0.0045s, with the total inference time of the 1D CNN models being 0.004s and the computation of the features taking 0.0005s.
III-B Evaluation Metrics
@@
The algorithm’s performance online, following the offline analysis in [ref:mungai2024fall], is evaluated using lead time, false positive rate, and response time. In addition to these metrics, we introduce a new metric to assess the algorithm’s effectiveness: the recovery rate. The recovery rate measures the number of unsafe trajectories that are stabilized by switching to a recovery controller when a critical fault is detected. We report the recovery rate only for simulation tests due to the difficulty of replicating the same fault multiple times on the hardware. Since implementing a recovery controller falls outside the scope of this work, we utilize Agility’s base controller, which already incorporates a fall prevention strategy. We note that only abrupt and incipient faults are considered in our evaluation, as the nominal fall prediction algorithm’s effectiveness in detecting intermittent faults, as demonstrated in [ref:mungai2024fall], depends on accurately identifying critical abrupt and incipient faults.
III-C Experimental Setup
@@
Abrupt and incipient faults are emulated online in simulation by applying impulsive and trapezoidal pushing forces, respectively, to regions F in Fig. 6. To assess the fall prediction algorithm’s performance online in simulation and compare it to the offline results presented in [ref:mungai2024fall], we replicate the 200 abrupt and incipient tests used in the offline evaluation and apply the algorithm to these tests. The simulation is run on a machine with an Intel i9-7900x processor.
In the hardware setup, we do not have direct access to the main computer, so we interact with the robot by sending commands and receiving state information over UDP at approximately 1 kHz. This information is processed by the nominal controller, which publishes topics via ROS also at roughly 1 kHz. The robot is connected to an Alienware m15 computer with an Intel i9-12900H processor via Ethernet. This computer runs both the fall prediction algorithm and the nominal controller.
We execute the algorithm online in hardware 13 times, emulating 8 abrupt faults and 6 incipient faults. In contrast to the simulation setup, faults are introduced by applying pushing forces using a pole to region C in Fig. 6, as region F houses the on-board computer. To capture the gradual nature of incipient faults, the pole is first positioned on the robot before applying force. To prevent potential damage during experiments, the robot is secured to a gantry using a slack cable. Note that the nominal fall prediction algorithm is transferred from hardware to simulation in a zero-shot manner - no additional data is used for training, and no modifications are made to the algorithm.
III-D Performance Analysis
III-D1 Simulation — Offline vs Online
@@
As shown in Table I, the true average lead time for the online execution in simulation is 0.01s higher than the offline execution. This minor discrepancy in fault detection is primarily due to communication delays, which cause the online algorithm to occasionally run slower, averaging 40 Hz but sometimes dipping to lower rates. As a result, windows with features are sampled at intervals slightly greater or less than 40 Hz, leading to differences in the timing of critical fault detection. These timing variations also contribute to small changes in the average response time and predicted lead time, which decrease by 0.02s and increase by 0.04s, respectively, compared to the offline results. The difference between the average true lead time and predicted lead time is 0.03s for the online execution and 0.02s for the offline execution. Both online and offline evaluations report an average false positive rate of 0. Given the small differences between the online and offline results, we can conclude that the algorithm’s online performance in simulation is closely aligned with its offline results. Furthermore, the 0.97 recovery rate for the unsafe trajectories demonstrates the effectiveness of running the fall prediction algorithm online. Fig. 4 showcases the results of switching vs not switching to a recovery controller when the nominal fall prediction algorithm detects a critical fault.
@@
TABLE I: Online vs Offline Average Simulation Results of the Nominal Fall Prediction Algorithm Evaluated on Abrupt and Incipient Faults
Operation Predicted Lead Time True Lead Time False Positive Rate Response Time (s) Recovery Rate
Offline[ref:mungai2024fall] 1.18 1.16 0.0 0.44 N/A
@@
III-D2 Hardware — Offline vs Online
In the hardware tests online, 4 of the 8 abrupt faults are detected as critical faults, and 3 of the 6 incipient faults are identified as critical faults. The average predicted lead time is 1.04s which is similar to the offline hardware results reported in [ref:mungai2024fall]. Note that we only report the average predicted lead time, as measuring the force application time and hence the response time online without force sensors proved challenging. Additionally, false-positive rates are unknown since the recovery controller is triggered whenever falls are detected. However, the similarity in average predicted lead times between the online and offline implementations indicates that the nominal fall prediction algorithm’s performance on hardware remains consistent with its offline results and the simulation results. Fig. 5 demonstrates the fall prediction algorithm detecting a critical fault online in hardware.
III-D3 Simulation vs Hardware — Online
The average predicted lead time achieved online in hardware is 0.1s lower than the results obtained in the online simulation. This deviation can be attributed to the smaller number of hardware online tests, which do not encompass as broad a range of critical faults. In contrast, the average runtime of the fall prediction algorithm for the hardware setup is 0.0061s, which is comparable to the algorithm’s runtime in simulation. The 0.0016s difference in runtime is negligible, given the algorithm’s sampling rate of 0.025s (40Hz). Furthermore, the nominal controller publishes at about 1kHz in both hardware and simulation. The negligible difference in runtime for the fall prediction algorithm between simulation and hardware, the small sim-to-real gap demonstrated offline in [ref:mungai2024fall], and the similarity in performance when executed offline and online in simulation, as shown in Table I, suggest that the algorithm’s performance offline in simulation reliably translates online to hardware, further validating its real-world applicability. Fig. 1 offers a consolidated view of critical fault detection in both hardware and simulation, with the recovery algorithm in the loop, as illustrated in Figs. 4 and 5.